# Artem Lopatin Artem Lopatin Universidade Estadual de Campinas (UNICAMP), 651 Sergio Buarque de Holanda, 13083-859 Campinas, SP, Brazil

### v2

@@ Abstract.
Keywords: polynomial equations, octonions, split-octonions, positive characteristic.
2020 MSC: 17-08, 17A75, 17D05, 20G41, 65H05.
1. Introduction
Unless stated otherwise, let $\mathbb{F}$ be a field (which may be finite) of arbitrary characteristic $\mathop{\rm char}\mathbb{F}\geq 0$. All vector spaces and algebras are assumed to be defined over $\mathbb{F}$.
1.1. Equations over octonions
@@ 2. $\bullet$
Flaut and Shpakivskyi [16] studied the equation $x^{n}=a$ over real octonion division algebras. For an octonion division algebra $\mathbf{C}$ over an arbitrary field $\mathbb{F}$, Chapman [14] developed a complete method for finding the solutions of a general polynomial equation $a_{n}x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}=0$ over $\mathbf{C}$. Furthermore, Chapman and Vishkautsan [12], working over a division algebra $\mathbf{C}$, determined the solutions of the polynomial equation $(a_{n}c)x^{n}+(a_{n-1}c)x^{n-1}+\cdots+(a_{1}c)x+(a_{0}c)=0$, and also discussed the solutions of the related equation $(ca_{n})x^{n}+(ca_{n-1})x^{n-1}+\cdots+(ca_{1})x+(ca_{0})=0$. Chapman and Levin [11] introduced a method for finding so-called alternating roots of polynomials over an arbitrary division Cayley–Dickson algebra. In a subsequent work, Chapman and Vishkautsan [13] investigated conditions under which, for a root $a$ of a polynomial $f(x)$ over a general Cayley–Dickson algebra, there exists a factorization f(x)=g(x)(xa)f(x)=g(x)(x-a) for some polynomial $g(x)$. Working over the split-octonion algebra over an algebraically closed field, Lopatin and Rybalov [25] solved all polynomial equations in which all coefficients except the constant term are scalar. As a consequence, the $n$-th roots of a split-octonion were computed. In [27], Lopatin and Zubkov studied the linear equations $ax=c$, $(ax)b=c$, and $a(bx)=c$ over the split-octonion algebra $\mathbf{O}$ when the base field $\mathbb{F}$ is algebraically closed. It is worth noting that, over a division octonion algebra, these equations are easily solvable and admit a unique solution whenever $a,b\neq 0$. In contrast, over an algebraically closed field the situation is substantially more delicate. As a consequence of the main result in [27], it was shown that if a linear monomial equation over octonions with a non-zero constant term has at least two solutions, then it necessarily admits an invertible solution.
@@
1.2. Results
In Section 2 we define the octonion algebra $\mathbf{O}$ and its automorphism group $\mathop{\rm Aut}(\mathbf{O})$. In Section 3 we extend the results of [25] from the case of an algebraically closed field to the case of an arbitrary field, which includes, in particular, the case of $\mathbb{R}$ that is important for applications in physics.
@@ 1.2. Results
(1.1) $\alpha_{n}x^{n}+\alpha_{n-1}x^{n-1}+\cdots+\alpha_{1}x=c$
with scalar coefficients $\alpha_{1},\ldots,\alpha_{n}\in\mathbb{F}$ and a possibly non-scalar constant term $c\in\mathbf{O}$, where the variable is $x\in\mathbf{O}$ (see Theorem 3.3). The solution of equation (1.1) is reduced to solving polynomial equations over the field $\mathbb{F}$. In Corollary 3.4 we provide further details for the case where $c\in\mathbf{O}$ is non-scalar. As applications, we consider
@@ 2. $\bullet$
In Corollary 3.5 we show that, when $c$ is non-scalar, the number of solutions of equation (1.1) is finite.
@@
2. Octonions
2.1. Split-octonions
@@
Similarly to $\mathbf{O}(\mathbb{F})$, we define the algebra of octonions $\mathbf{O}(\mathcal{A})$ over any commutative associative $\mathbb{F}$-algebra $\mathcal{A}$.
@@
The linear trace function is given by $\mathop{\rm tr}(a)=\alpha+\beta$. Note that
$\mathop{\rm tr}(a)\,1_{\mathbf{O}}=a+\overline{a}\qquad\text{and}\qquad n(a)\,1_{\mathbf{O}}=a\,\overline{a}.$
@@
$\mathbf{O}_{0}=\{a\in\mathbf{O}\mid\mathop{\rm tr}(a)=0\},$
@@
(2.1) $\mathop{\rm tr}(aa^{\prime})=\mathop{\rm tr}(a^{\prime}a),\qquad n(aa^{\prime})=n(a)\,n(a^{\prime}),$
@@
(2.2) $n(a+a^{\prime})=n(a)+n(a^{\prime})-\mathop{\rm tr}(aa^{\prime})+\mathop{\rm tr}(a)\mathop{\rm tr}(a^{\prime}).$
@@
(2.5) $\overline{a}(ab)=n(a)\,b,\qquad(ba)\overline{a}=n(a)\,b.$
@@ 2. $\bullet$
As a direct consequence of identities (2.5), for every $a\in\mathbf{O}\setminus\mathbf{O}_{\#}$ we have
(2.6) $a^{-1}(ab)=b,\qquad(ba)a^{-1}=b,$
@@
2.2. The group $\mathop{\rm Aut}(\mathbf{O})$
The group $\mathop{\rm Aut}(\mathbf{O})$ of all automorphisms of the algebra $\mathbf{O}$ is the exceptional simple algebraic group ${\rm G}_{2}(\mathbb{F})$ when the field $\mathbb{F}$ is algebraically closed. In the general case, the group $\mathop{\rm Aut}(\mathbf{O})$ contains a subgroup isomorphic to ${\rm SL}_{3}(\mathbb{F})$. More precisely, every element $g\in{\rm SL}_{3}(\mathbb{F})$ defines an automorphism of $\mathbf{O}$ by
@@ 2.2. The group $\mathop{\rm Aut}(\mathbf{O})$
where $g^{-T}$ denotes $(g^{-1})^{T}$ and the vectors $\mathbf{u},\mathbf{v}\in\mathbb{F}^{3}$ are regarded as row vectors. For every $\mathbf{u},\mathbf{v}\in\mathbb{F}^{3}$, define automorphisms $\delta_{1}(\mathbf{u}),\delta_{2}(\mathbf{v})\in\mathop{\rm Aut}(\mathbf{O})$ by
@@
(2.7) $\hbar\colon\mathbf{O}\to\mathbf{O},\qquad a\longmapsto\left(\begin{array}[]{cc}\beta&-\mathbf{v}\\ -\mathbf{u}&\alpha\\ \end{array}\right),$
also belongs to $\mathop{\rm Aut}(\mathbf{O})$.
The action of $\mathop{\rm Aut}(\mathbf{O})$ on $\mathbf{O}$ satisfies the following properties:
(2.8) $\overline{ga}=g\overline{a},\qquad\mathop{\rm tr}(ga)=\mathop{\rm tr}(a),\qquad n(ga)=n(a),\qquad q(ga,ga^{\prime})=q(a,a^{\prime}),$
for all $g\in\mathop{\rm Aut}(\mathbf{O})$ and $a,a^{\prime}\in\mathbf{O}$ (see, for example, equation (2.2) in [1]). Consequently, the group $\mathop{\rm Aut}(\mathbf{O})$ also acts naturally on the subspaces $\mathbf{O}_{0}$ and $\mathbf{O}_{\#}$.
2.3. Canonical octonions
A minimal set of representatives for the $\mathop{\rm Aut}(\mathbf{O})$-orbits in $\mathbf{O}$ was described in Proposition 3.3 of [26] in the case where the base field is algebraically closed. We extend this result to an arbitrary field in the following proposition. The elements appearing in Proposition 2.2 will be referred to as canonical octonions.
@@ 2.3. Canonical octonions
A minimal set of representatives for the $\mathop{\rm Aut}(\mathbf{O})$-orbits in $\mathbf{O}$ consists of the following elements:
@@
where $\alpha,\beta\in\mathbb{F}$. In other words, $\mathbf{O}$ is the disjoint union of the following $\mathop{\rm Aut}(\mathbf{O})$-orbits:
$\alpha\,1_{\mathbf{O}}\qquad\text{and}\qquad O(\alpha,\beta):=\mathop{\rm Aut}(\mathbf{O})\cdot\left(\begin{array}[]{cc}\alpha&(\beta,0,0)\\ (1,0,0)&0\\ \end{array}\right),$
@@
Proof.
By Lemma 3.3 of [1], two elements $a,b\in\mathbf{O}\setminus\mathbb{F}1_{\mathbf{O}}$ belong to the same $\mathop{\rm Aut}(\mathbf{O})$-orbit if and only if $\mathop{\rm tr}(a)=\mathop{\rm tr}(b)$ and $n(a)=n(b)$. The proof is completed by observing that
$\mathop{\rm tr}\!\left(\left(\begin{array}[]{cc}\alpha&(\beta,0,0)\\ (1,0,0)&0\\ \end{array}\right)\right)=\alpha,\qquad n\!\left(\left(\begin{array}[]{cc}\alpha&(\beta,0,0)\\ (1,0,0)&0\\ \end{array}\right)\right)=-\beta.$
The above proof of Proposition 2.2 is non-constructive and relies on Lemma 3.3 of [1]. For the sake of completeness, we present an alternative proof that does not depend on Lemma 3.3 of [1] and that, for every $a\in\mathbf{O}$, explicitly constructs an element $g\in\mathop{\rm Aut}(\mathbf{O})$ such that $ga$ is in canonical form. This constructive approach may be useful for practical computations involving octonions.
Constructive proof of Proposition 2.2. In this proof, we use the symbol $\ast$ to denote an arbitrary element of $\mathbb{F}$. Let $a=\left(\begin{array}[]{cc}\alpha_{1}&\mathbf{u}\\ \mathbf{v}&\alpha_{8}\\ \end{array}\right)$. Then one of the following two cases occurs.
@@
If $\alpha_{8}=0$, then $a$ is canonical. Otherwise, acting on $a$ with $\delta_{1}(-\alpha_{8},0,0)$, we obtain the canonical octonion $\left(\begin{array}[]{cc}\alpha_{1}+\alpha_{8}&(-\alpha_{1}\alpha_{8},0,0)\\ (1,0,0)&\alpha_{8}\\ \end{array}\right)$.
@@
As above, using part (a) of Lemma 3.2 from [26] together with the automorphism $\hbar$, we further reduce to the octonion
@@ 2. $\bullet$
    $\mathbf{u}^{\prime\prime}=(0,1,0)$. If $\alpha\neq 0$, we apply $\delta_{1}\!\left(0,-\frac{1}{\alpha},0\right)$ to obtain the canonical octonion $\left(\begin{array}[]{cc}\alpha&\mathbf{0}\\ (1,0,0)&0\\ \end{array}\right)$. If $\alpha=0$, we apply $\delta_{2}(0,0,1)$ to obtain the canonical octonion $\left(\begin{array}[]{cc}0&\mathbf{0}\\ (1,0,0)&0\\ \end{array}\right)$.
$\Box$
3. Polynomial equations
@@
Given $a\in\mathbf{O}$, we denote $\mathop{\rm tr}(a)=\alpha$ and $n(a)=-\beta$, where $\alpha,\beta\in\mathbb{F}$.
@@
    (3.1) $a^{n}=p_{n-1}(\alpha,\beta)\,a+p_{n-2}(\alpha,\beta)\,\beta 1_{\mathbf{O}}.$
@@
    (3.2) $f(a)=\widehat{f}(\alpha,\beta)\,a+\widecheck{f}(\alpha,\beta)\,\beta 1_{\mathbf{O}}.$
Proof.
1. For brevity, set $q_{n}=p_{n}(\alpha,\beta)$. The quadratic equation (2.3) can be rewritten in the form
(3.3) $a^{2}=\alpha a+\beta 1_{\mathbf{O}}.$
We prove formula (3.1) by induction on $n\geqslant 1$. The case $n=1$ is trivial.
Assume that (3.1) holds for $n=k$. Then, by the induction hypothesis and (3.3), we obtain
@@
    Assume $c=\gamma 1_{\mathbf{O}}$ for some $\gamma\in\mathbb{F}$. Then X ={ν1_O|  νF satisfies f(ν)=γ}{ O(λ,μ)  |  λ,μF satisfy ^f(λ,μ)=0 and μ ˇf(λ,μ)=γ}.
@@
    (3.4) $\left\{\begin{array}[]{rcl}\lambda\,\widehat{f}(\lambda,\mu)+2\mu\,\widecheck{f}(\lambda,\mu)&=&\mathop{\rm tr}(c),\\[2.84526pt] -\mu\,\widehat{f}(\lambda,\mu)^{2}+\lambda\mu\,\widehat{f}(\lambda,\mu)\widecheck{f}(\lambda,\mu)+\mu^{2}\,\widecheck{f}(\lambda,\mu)^{2}&=&n(c),\\[2.84526pt] \widehat{f}(\lambda,\mu)&\neq&0,\end{array}\right$.
@@
    (3.5) $x=\frac{1}{\widehat{f}(\lambda,\mu)}\Big(c-\mu\,\widecheck{f}(\lambda,\mu)1_{\mathbf{O}}\Big).$
    In this case, we have $\mathop{\rm tr}(x)=\lambda$ and $n(x)=-\mu$.
Proof.
1. Let $c=\gamma 1_{\mathbf{O}}$ for some $\gamma\in\mathbb{F}$, and consider an arbitrary $x\in\mathbf{O}$. By Proposition 2.2, there exists $g\in\mathop{\rm Aut}(\mathbf{O})$ such that $gx$ is canonical. Note that the equality $f(x)=c$ is equivalent to
(3.6) $f(gx)=\gamma 1_{\mathbf{O}}.$
@@
    $gx=\nu 1_{\mathbf{O}}$ for some $\nu\in\mathbb{F}$. In this case, equality (3.6) is equivalent to $f(\nu)=\gamma$. Hence, x = g^-1(ν1_O) = ν1_O.
@@
    $gx=\left(\begin{array}[]{cc}\lambda&(\mu,0,0)\\ (1,0,0)&0\\ \end{array}\right)$ for some $\lambda,\mu\in\mathbb{F}$. By part 2 of Proposition 3.1, equality (3.6) is equivalent to { ^f(λ,μ)=0,μ ˇf(λ,μ)=γ. Since $x\in O(\lambda,\mu)$, we have that $x\in X$ if and only if $O(\lambda,\mu)\subseteq X$.
@@
(3.7) $\widehat{f}(\lambda,\mu)\,x+\mu\,\widecheck{f}(\lambda,\mu)\,1_{\mathbf{O}}=c,$
where $\lambda=\mathop{\rm tr}(x)$ and $\mu=-n(x)$. If $\widehat{f}(\lambda,\mu)=0$, then equality (3.7) would imply $c\in\mathbb{F}1_{\mathbf{O}}$, a contradiction. Hence, $x$ is given by equality (3.5).
Applying the trace and norm to equality (3.7) and using the linearity of the trace together with formula (2.2), we obtain that system (3.4) holds.
(b) Conversely, assume that $x\in\mathbf{O}$ is defined by equality (3.5) for some $\lambda,\mu\in\mathbb{F}$ satisfying system (3.4). System (3.4) implies
$\mathop{\rm tr}(x)=\frac{1}{\widehat{f}(\lambda,\mu)}\Big(\mathop{\rm tr}(c)-2\mu\,\widecheck{f}(\lambda,\mu)\Big)=\lambda.$
Similarly, formula (2.2) together with system (3.4) gives
$n(x)=\frac{1}{\widehat{f}(\lambda,\mu)^{2}}\Big(n(c)+\mu^{2}\widecheck{f}(\lambda,\mu)^{2}-\mu\,\mathop{\rm tr}(c)\widecheck{f}(\lambda,\mu)\Big)\\ =\frac{1}{\widehat{f}(\lambda,\mu)^{2}}\Big(n(c)-\lambda\mu\,\widehat{f}(\lambda,\mu)\widecheck{f}(\lambda,\mu)-\mu^{2}\widecheck{f}(\lambda,\mu)^{2}\Big)=-\mu.$
Therefore, by part 2 of Proposition 3.1 and equality (3.5),
@@
    Assume $\mathop{\rm char}\mathbb{F}\neq 2$. Then $x\in X$ if and only if there exist $\lambda,\mu\in\mathbb{F}$ satisfying
    (3.8) $\left\{\begin{array}[]{rcl}(\lambda^{2}+4\mu)\widehat{f}(\lambda,\mu)^{2}&=&\mathop{\rm tr}(c)^{2}-4\,n(c),\\[2.84526pt] \lambda\widehat{f}(\lambda,\mu)+2\mu\widecheck{f}(\lambda,\mu)&=&\mathop{\rm tr}(c),\\[2.84526pt] \widehat{f}(\lambda,\mu)&\neq&0,\end{array}\right$.
@@
    (3.9) $x=\frac{1}{\widehat{f}(\lambda,\mu)}\Big(c-\frac{\mathop{\rm tr}(c)}{2}1_{\mathbf{O}}\Big)+\frac{\lambda}{2}\,1_{\mathbf{O}}.$
@@
    Assume $\mathop{\rm char}\mathbb{F}=2$.

### v4

# On polynomial equations over split-octonions: the arbitrary field case

@@ Abstract
keywords:
Polynomial equations, Octonions, Split-octonions, Positive characteristic.
[…]
Acknowledgements.
1 Introduction
Unless stated otherwise, let $\mathbb{F}$ be a field (which may be finite) of arbitrary characteristic $\operatorname{char}\mathbb{F}\geq 0$. All vector spaces and algebras are assumed to be defined over $\mathbb{F}$.
1.1 Equations over octonions
@@ 2. $\bullet$
Flaut and Shpakivskyi [16] studied the equation $x^{n}=a$ over real octonion division algebras. For an octonion division algebra $\mathbf{C}$ over an arbitrary field $\mathbb{F}$, Chapman [11] developed a complete method for finding the solutions of a general polynomial equation $a_{n}x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0}=0$ over $\mathbf{C}$. Furthermore, Chapman and Vishkautsan [13], working over a division algebra $\mathbf{C}$, determined the solutions of the polynomial equation $(a_{n}c)x^{n}+(a_{n-1}c)x^{n-1}+\cdots+(a_{1}c)x+(a_{0}c)=0$, and also discussed the solutions of the related equation $(ca_{n})x^{n}+(ca_{n-1})x^{n-1}+\cdots+(ca_{1})x+(ca_{0})=0$. Chapman and Levin [12] introduced a method for finding so-called alternating roots of polynomials over an arbitrary division Cayley–Dickson algebra. In a subsequent work, Chapman and Vishkautsan [14] investigated conditions under which, for a root $a$ of a polynomial $f(x)$ over a general Cayley–Dickson algebra, there exists a factorization f(x)=g(x)(xa)f(x)=g(x)(x-a) for some polynomial $g(x)$. Working over the split-octonion algebra over an algebraically closed field, Lopatin and Rybalov [25] solved all polynomial equations in which all coefficients except the constant term are scalar. As a consequence, the $n$-th roots of a split-octonion were computed. In [28], Lopatin and Zubkov studied the linear equations given by $ax=c$, $(ax)b=c$ and $a(bx)=c$ over the split-octonion algebra $\mathbf{O}$ when the base field $\mathbb{F}$ is algebraically closed. It is worth noting that, over a division octonion algebra, these equations are easily solvable and admit a unique solution whenever $a,b\neq 0$. In contrast, over an algebraically closed field the situation is substantially more delicate. As a consequence of the main result in [28], it was shown that if a linear monomial equation over octonions with a non-zero constant term has at least two solutions, then it necessarily admits an invertible solution.
@@
1.2 Results
In Section 2 we define the octonion algebra $\mathbf{O}$ and its automorphism group $\operatorname{Aut}(\mathbf{O})$. In Section 3 we extend the results of [25] from the case of an algebraically closed field to the case of an arbitrary field, which includes, in particular, the case of $\mathbb{R}$ that is important for applications in physics.
@@ 1.2 Results
$\alpha_{n}x^{n}+\alpha_{n-1}x^{n-1}+\cdots+\alpha_{1}x=c$ (1)
with scalar coefficients $\alpha_{1},\ldots,\alpha_{n}\in\mathbb{F}$ and a possibly non-scalar constant term $c\in\mathbf{O}$, where the variable is $x\in\mathbf{O}$ (see Theorem 3.4). The solution of equation (1) is reduced to solving polynomial equations over the field $\mathbb{F}$. In Corollary 3.6 we provide further details for the case where $c\in\mathbf{O}$ is non-scalar. As applications, we consider
@@ 2. $\bullet$
In Corollary 3.7 we show that, when $c$ is non-scalar, the number of solutions of equation (1) is finite.
@@
2 Octonions
The definitions presented in this section are taken from [26]. Additional material on octonions can be found in the books [34, 29].
2.1 Split-octonions
@@
The linear trace function is given by $\operatorname{tr}(a)=\alpha+\beta$. Note that
$\operatorname{tr}(a)\,1_{\mathbf{O}}=a+\overline{a}\qquad\text{and}\qquad n(a)\,1_{\mathbf{O}}=a\,\overline{a}.$
@@
$\mathbf{O}_{0}=\{a\in\mathbf{O}\mid\operatorname{tr}(a)=0\},$
@@
$\operatorname{tr}(aa^{\prime})=\operatorname{tr}(a^{\prime}a),\qquad n(aa^{\prime})=n(a)\,n(a^{\prime}),$ (2)
@@
$n(a+a^{\prime})=n(a)+n(a^{\prime})-\operatorname{tr}(aa^{\prime})+\operatorname{tr}(a)\operatorname{tr}(a^{\prime}).$ (3)
@@
$\overline{a}(ab)=n(a)\,b,\qquad(ba)\overline{a}=n(a)\,b.$ (6)
@@ 2. $\bullet$
As a direct consequence of identities (6), for every $a\in\mathbf{O}\setminus\mathbf{O}_{\#}$ we have
$a^{-1}(ab)=b,\qquad(ba)a^{-1}=b,$ (7)
@@
2.2 The group $\operatorname{Aut}(\mathbf{O})$
The group $\operatorname{Aut}(\mathbf{O})$ of all automorphisms of the algebra $\mathbf{O}$ is the exceptional simple algebraic group ${\rm G}_{2}(\mathbb{F})$ when the field $\mathbb{F}$ is algebraically closed. In the general case, the group $\operatorname{Aut}(\mathbf{O})$ contains a subgroup isomorphic to $\operatorname{SL}_{3}(\mathbb{F})$. More precisely, every element $g\in\operatorname{SL}_{3}(\mathbb{F})$ defines an automorphism of $\mathbf{O}$ by
@@ 2.2 The group $\operatorname{Aut}(\mathbf{O})$
where $g^{-T}$ denotes $(g^{-1})^{T}$ and the vectors $\mathbf{u},\mathbf{v}\in\mathbb{F}^{3}$ are regarded as row vectors. For every $\mathbf{u},\mathbf{v}\in\mathbb{F}^{3}$, define automorphisms $\delta_{1}(\mathbf{u}),\delta_{2}(\mathbf{v})\in\operatorname{Aut}(\mathbf{O})$ by
@@
$\hbar\colon\mathbf{O}\to\mathbf{O},\qquad a\longmapsto\left(\begin{array}[]{cc}\beta&-\mathbf{v}\\ -\mathbf{u}&\alpha\\ \end{array}\right),$ (8)
also belongs to $\operatorname{Aut}(\mathbf{O})$.
The action of $\operatorname{Aut}(\mathbf{O})$ on $\mathbf{O}$ satisfies the following properties:
$\overline{ga}=g\overline{a},\qquad\operatorname{tr}(ga)=\operatorname{tr}(a),\qquad n(ga)=n(a),\qquad q(ga,ga^{\prime})=q(a,a^{\prime}),$ (9)
for all $g\in\operatorname{Aut}(\mathbf{O})$ and $a,a^{\prime}\in\mathbf{O}$ (see, for example, equation (2.2) in [1]). Consequently, the group $\operatorname{Aut}(\mathbf{O})$ also acts naturally on the subspaces $\mathbf{O}_{0}$ and $\mathbf{O}_{\#}$.
2.3 Canonical octonions
A minimal set of representatives for the $\operatorname{Aut}(\mathbf{O})$-orbits in $\mathbf{O}$ was described in Proposition 3.3 of [27] in the case where the base field is algebraically closed. We extend this result to an arbitrary field in the following proposition. The elements appearing in Proposition 2.2 will be referred to as canonical octonions.
@@ 2.3 Canonical octonions
A minimal set of representatives for the $\operatorname{Aut}(\mathbf{O})$-orbits in $\mathbf{O}$ consists of the following elements:
@@
where $\alpha,\beta\in\mathbb{F}$. In other words, $\mathbf{O}$ is the disjoint union of the following $\operatorname{Aut}(\mathbf{O})$-orbits:
$\alpha\,1_{\mathbf{O}}\qquad\text{and}\qquad O(\alpha,\beta):=\operatorname{Aut}(\mathbf{O})\cdot\left(\begin{array}[]{cc}\alpha&(\beta,0,0)\\ (1,0,0)&0\\ \end{array}\right),$
@@
Proof 2.3.
By Lemma 3.3 of [1], two elements $a,b\in\mathbf{O}\setminus\mathbb{F}1_{\mathbf{O}}$ belong to the same $\operatorname{Aut}(\mathbf{O})$-orbit if and only if $\operatorname{tr}(a)=\operatorname{tr}(b)$ and $n(a)=n(b)$. The proof is completed by observing that
$\operatorname{tr}\!\left(\left(\begin{array}[]{cc}\alpha&(\beta,0,0)\\ (1,0,0)&0\\ \end{array}\right)\right)=\alpha,\qquad n\!\left(\left(\begin{array}[]{cc}\alpha&(\beta,0,0)\\ (1,0,0)&0\\ \end{array}\right)\right)=-\beta.$
The above proof of Proposition 2.2 is non-constructive and relies on Lemma 3.3 of [1]. For the sake of completeness, we present an alternative proof that does not depend on Lemma 3.3 of [1] and that, for every $a\in\mathbf{O}$, explicitly constructs an element $g\in\operatorname{Aut}(\mathbf{O})$ such that $ga$ is in canonical form. This constructive approach may be useful for practical computations involving octonions.
Proof 2.4 (Constructive proof of Proposition 2.2).
In this proof, we use the symbol $\ast$ to denote an arbitrary element of $\mathbb{F}$. Let $a=\left(\begin{array}[]{cc}\alpha_{1}&\mathbf{u}\\ \mathbf{v}&\alpha_{8}\\ \end{array}\right)$. Then one of the following two cases occurs.
@@
If $\alpha_{8}=0$, then $a$ is canonical. Otherwise, acting on $a$ with $\delta_{1}(-\alpha_{8},0,0)$, we obtain the canonical octonion $\left(\begin{array}[]{cc}\alpha_{1}+\alpha_{8}&(-\alpha_{1}\alpha_{8},0,0)\\ (1,0,0)&0\\ \end{array}\right)$.
@@
As above, similarly to part (a) of Lemma 3.2 from [27], we further reduce to the octonion
@@ 2. $\bullet$
    $\mathbf{u}^{\prime\prime}=(0,1,0)$. If $\alpha\neq 0$, we apply $\delta_{1}\!\left(0,-\frac{1}{\alpha},0\right)$ to obtain the canonical octonion $\left(\begin{array}[]{cc}\alpha&\mathbf{0}\\ (1,0,0)&0\\ \end{array}\right)$. If $\alpha=0$, we apply $\delta_{2}(0,0,1)$ to obtain the canonical octonion $\left(\begin{array}[]{cc}0&\mathbf{0}\\ (1,0,0)&0\\ \end{array}\right)$,
and the proof is finished.
3 Polynomial equations
@@
Given $a\in\mathbf{O}$, we denote $\operatorname{tr}(a)=\alpha$ and $n(a)=-\beta$, where $\alpha,\beta\in\mathbb{F}$.
@@
    $a^{n}=p_{n-1}(\alpha,\beta)\,a+p_{n-2}(\alpha,\beta)\,\beta 1_{\mathbf{O}}.$ (10)
@@
    $f(a)=\widehat{f}(\alpha,\beta)\,a+\widecheck{f}(\alpha,\beta)\,\beta 1_{\mathbf{O}}.$ (11)
Proof 3.2.
1. For brevity, set $q_{n}=p_{n}(\alpha,\beta)$. The quadratic equation (4) can be rewritten in the form
$a^{2}=\alpha a+\beta 1_{\mathbf{O}}.$ (12)
We prove formula (10) by induction on $n\geqslant 1$. The case $n=1$ is trivial.
Assume that (10) holds for $n=k$. Then, by the induction hypothesis and (12), we obtain
@@
Recall that $O(\alpha,\beta)$ is defined in Proposition 2.2 as the $\operatorname{Aut}(\mathbf{O})$-orbit of the canonical octonion $\left(\begin{array}[]{cc}\alpha&(\beta,0,0)\\ (1,0,0)&0\\ \end{array}\right)$.
@@
    Assume $c=\gamma 1_{\mathbf{O}}$ for some $\gamma\in\mathbb{F}$. Then
    $X=\big\{\nu 1_{\mathbf{O}}\,\big|\,\nu\in\mathbb{F}\text{ satisfies }f(\nu)=\gamma\big\}\;\cup\bigcup_{\begin{subarray}{c}\lambda,\mu\in\mathbb{F}\\ \widehat{f}(\lambda,\mu)=0\\ \mu\,\widecheck{f}(\lambda,\mu)=\gamma\end{subarray}}O(\lambda,\mu).$
@@
    $\left\{\begin{array}[]{rcl}\lambda\,\widehat{f}(\lambda,\mu)+2\mu\,\widecheck{f}(\lambda,\mu)&=&\operatorname{tr}(c),\\[2.84526pt] -\mu\,\widehat{f}(\lambda,\mu)^{2}+\lambda\mu\,\widehat{f}(\lambda,\mu)\widecheck{f}(\lambda,\mu)+\mu^{2}\,\widecheck{f}(\lambda,\mu)^{2}&=&n(c),\\[2.84526pt] \widehat{f}(\lambda,\mu)&\neq&0,\end{array}\right$. (13)
@@
    $x=\frac{1}{\widehat{f}(\lambda,\mu)}\Big(c-\mu\,\widecheck{f}(\lambda,\mu)1_{\mathbf{O}}\Big).$ (14)
    In this case, we have $\operatorname{tr}(x)=\lambda$ and $n(x)=-\mu$.
Proof 3.5.
1. Let $c=\gamma 1_{\mathbf{O}}$ for some $\gamma\in\mathbb{F}$, and consider an arbitrary $x\in\mathbf{O}$. By Proposition 2.2, there exists $g\in\operatorname{Aut}(\mathbf{O})$ such that $gx$ is canonical. Note that the equality $f(x)=c$ is equivalent to
$f(gx)=\gamma 1_{\mathbf{O}}.$ (15)
@@
    $gx=\nu 1_{\mathbf{O}}$ for some $\nu\in\mathbb{F}$. In this case, equality (15) is equivalent to $f(\nu)=\gamma$. Hence,
    $x=g^{-1}(\nu 1_{\mathbf{O}})=\nu 1_{\mathbf{O}}.$
@@
    $gx=\left(\begin{array}[]{cc}\lambda&(\mu,0,0)\\ (1,0,0)&0\\ \end{array}\right)$ for some $\lambda,\mu\in\mathbb{F}$. By part 2 of Proposition 3.1, equality (15) is equivalent to
[…]
    Since $x\in O(\lambda,\mu)$, we have that $x\in X$ if and only if $O(\lambda,\mu)\subseteq X$.
@@
$\widehat{f}(\lambda,\mu)\,x+\mu\,\widecheck{f}(\lambda,\mu)\,1_{\mathbf{O}}=c,$ (16)
where $\lambda=\operatorname{tr}(x)$ and $\mu=-n(x)$. If $\widehat{f}(\lambda,\mu)=0$, then equality (16) would imply $c\in\mathbb{F}1_{\mathbf{O}}$, a contradiction. Hence, $x$ is given by equality (14).
Applying the trace and norm to equality (16) and using the linearity of the trace together with formula (3), we obtain that system (13) holds.
(b) Conversely, assume that $x\in\mathbf{O}$ is defined by equality (14) for some $\lambda,\mu\in\mathbb{F}$ satisfying system (13). System (13) implies
$\operatorname{tr}(x)=\frac{1}{\widehat{f}(\lambda,\mu)}\Big(\operatorname{tr}(c)-2\mu\,\widecheck{f}(\lambda,\mu)\Big)=\lambda.$
Similarly, formula (3) together with system (13) gives
$n(x)=\frac{1}{\widehat{f}(\lambda,\mu)^{2}}\Big(n(c)+\mu^{2}\widecheck{f}(\lambda,\mu)^{2}-\mu\,\operatorname{tr}(c)\widecheck{f}(\lambda,\mu)\Big)\\ =\frac{1}{\widehat{f}(\lambda,\mu)^{2}}\Big(n(c)-\lambda\mu\,\widehat{f}(\lambda,\mu)\widecheck{f}(\lambda,\mu)-\mu^{2}\widecheck{f}(\lambda,\mu)^{2}\Big)=-\mu.$
Therefore, by part 2 of Proposition 3.1 and equality (14),
@@
    Assume $\operatorname{char}\mathbb{F}\neq 2$. Then $x\in X$ if and only if there exist $\lambda,\mu\in\mathbb{F}$ satisfying
    $\left\{\begin{array}[]{rcl}(\lambda^{2}+4\mu)\widehat{f}(\lambda,\mu)^{2}&=&\operatorname{tr}(c)^{2}-4\,n(c),\\[2.84526pt] \lambda\widehat{f}(\lambda,\mu)+2\mu\widecheck{f}(\lambda,\mu)&=&\operatorname{tr}(c),\\[2.84526pt] \widehat{f}(\lambda,\mu)&\neq&0,\end{array}\right$. (17)
@@
    $x=\frac{1}{\widehat{f}(\lambda,\mu)}\Big(c-\frac{\operatorname{tr}(c)}{2}1_{\mathbf{O}}\Big)+\frac{\lambda}{2}\,1_{\mathbf{O}}.$ (18)
@@
    Assume $\operatorname{char}\mathbb{F}=2$.